### v1

# Position: Beyond Euclidean - Foundation Models Should Embrace Non-Euclidean Geometries

Abstract
In the era of foundation models and Large Language Models (LLMs), Euclidean space has been the de facto geometric setting for machine learning architectures. However, recent literature has demonstrated that this choice comes with fundamental limitations. At a large scale, real-world data often exhibit inherently non-Euclidean structures, such as multi-way relationships, hierarchies, symmetries, and non-isotropic scaling, in a variety of domains, such as languages, vision, and the natural sciences. It is challenging to effectively capture these structures within the constraints of Euclidean spaces. This position paper argues that moving beyond Euclidean geometry is not merely an optional enhancement but a necessity to maintain the scaling law for the next-generation of foundation models. By adopting these geometries, foundation models could more efficiently leverage the aforementioned structures. Task-aware adaptability that dynamically reconfigures embeddings to match the geometry of downstream applications could further enhance efficiency and expressivity. Our position is supported by a series of theoretical and empirical investigations of prevalent foundation models. Finally, we outline a roadmap for integrating non-Euclidean geometries into foundation models, including strategies for building geometric foundation models via fine-tuning, training from scratch, and hybrid approaches.
Machine Learning, ICML
Machine Learning, ICML
1 Introduction
Foundation models, such as Large Language Models (LLMs), have emerged as a cornerstone of current AI advancements, due to their ability to generalize across diverse tasks with minimal fine-tuning (Bommasani et al., 2021; Devlin et al., 2018; Brown et al., 2020; Ramesh et al., 2021). To date, Euclidean geometry has remained the default framework for designing such models. This preference is largely driven by the natural compatibility of Euclidean geometry with fundamental neural network operations—such as linear transformations, convolutions, and attention mechanisms—all of which can be efficiently executed using standard linear algebra in Euclidean space.
However, real-world datasets often exhibit implicit non-Euclidean structures. These structures manifest in many forms, such as the hierarchical organization of natural language—including concept taxonomies and entailment relationships (Nickel & Kiela, 2017; Tifrea et al., 2019; Le et al., 2019)—as well as hierarchical relationships among object classes, scenes, and their constituent categories in visual data (Ge et al., 2023; Pal et al., 2024). Furthermore, non-Euclidean characteristics are inherent in biological data, such as protein structures (Villegas-Morcillo et al., 2021) and RNA-seq data (Klimovskaia et al., 2020). Given these intrinsically non-Euclidean characteristics of the training data, combined with the ongoing challenges faced by foundation models—from hallucinations to computational inefficiencies—it becomes crucial to question whether Euclidean geometry remain the default for foundational models.
Position: This position paper argues that the development of non-Euclidean foundation models is essential for effectively representing, modeling, and analyzing complex data structures and relationships in real-world applications. With arguments grounded in both theoretical insights and experimental evidence, we contend that by aligning foundation models—whether visual, linguistic, or scientific—with the intrinsic geometries of their training data, we can improve three critical aspects of these models: representational capabilities, adaptability to diverse geometric structures in training datasets, and scalability.
Figure 1: Manifolds with corresponding graph structures or underlying relationships, which represent different types of token relationships: hierarchical (left), uniform (middle), and cyclical (right) dependencies.
Representational Capabilities. Euclidean space has limited capabilities to represent complex geometric structures with diverse local properties, resulting in significant distortion when embedding such data in low-dimensional Euclidean spaces (Tenenbaum et al., 2000). In contrast, hyperbolic spaces, with their negative curvature, excel at representing hierarchical and tree-like structures with minimal distortion in low dimensions (Krioukov et al., 2010; Sarkar, 2011). Similarly, spherical geometries, defined by positive curvature, are well-suited for modeling data with bounded structures and angular relationships (Durrant & Leontidis, 2022; Esteves et al., 2017; Trosten et al., 2023).
Adaptability. Incorporating non-Euclidean geometrical operations into foundation models provides substantially enhanced adaptability to the diverse geometric structures in training datasets, particularly in the case of large-scale datasets—as is typical for these models—where heterogeneity is inherent by design. This adaptability improves the models’ flexibility and transferability, as many instances of downstream tasks benefit from acknowledging the geometric structure in the data, such as knowledge graph completion (Bai et al., 2021; Wang et al., 2021), social network analysis (Yang et al., 2022d; Kennedy et al., 2013), multi-label classification, drug discovery (Poleksic, 2023), user preferences recommendation (Chen et al., 2022b; Yang et al., 2022c, b), and code understanding (Tang et al., 2023).
Low Dimensionality and Scalability. Beyond enhancing representation capabilities, adapting foundation models to the non-Euclidean geometry inherent in their training corpora enables lower-dimensional embeddings, reducing computational costs without sacrificing performance. A critical implication lies in the scaling laws of foundation models (Hoffmann et al., 2022), where performance in Euclidean models follows a power-law scaling of the form $L(N)\propto N^{-\alpha}$, with $L$ and $N$ being the loss and the number of parameters respectively. This scaling behavior reveals inherent inefficiencies in how Euclidean geometries handle increasing model complexity and data dimensionality. In contrast, adopting Riemannian manifolds with tailored curvature properties improves scaling by efficiently compressing information (Brehmer et al., 2024; Kiani et al., 2024). This improvement arises from the manifold’s ability to naturally capture long-range dependencies and hierarchical structures (Tifrea et al., 2019), which the Euclidean space struggles to represent effectively. Moreover, product manifolds and mixed-geometry approaches (Gu et al., 2019; Wang et al., 2021) further improve efficiency and performance by allowing different model components to scale according to their optimal geometric properties.
Moreover, we propose a roadmap for integrating non-Euclidean geometries into foundation models. This includes adapting existing Euclidean models to incorporate these principles and developing foundation models from scratch. We also highlight key challenges and outline the steps required to advance this non-Euclidean vision for foundation models.
2 Background and Preliminaries
In this section, we give an overview of non-Euclidean spaces, particularly focusing on Riemannian manifolds. For more details please see (Lee, 2013) and Appendix A.
2.1 Non-Euclidean Geometry Foundations
Riemannian Manifolds. A smooth $n$-dimensional manifold $\mathcal{M}$ is a topological space that is locally Euclidean. Each point $x$ is associated with a tangent space $T_{x}\mathcal{M}$, which is an $n$-dimensional vector space that acts as a first-order local approximation of $\mathcal{M}$. A Riemannian metric $\mathfrak{g}$ on $\mathcal{M}$ is a collection $\mathfrak{g}\coloneqq(\mathfrak{g}_{x})_{x\in\mathcal{M}}$ of positive definite bilinear forms $\mathfrak{g}_{x}(\cdot,\cdot):T_{x}\mathcal{M}\times T_{x}\mathcal{M}\to\mathbb{R}^{n}$, varying smoothly with $x$. $\mathfrak{g}_{x}$ induces the (sectional) curvature at point $x$, which measures how $\mathcal{M}$ deviates from flatness at $x$. A Riemannian manifold is a pairing $(\mathcal{M},\mathfrak{g})$. For example, $\mathbb{R}^{n}$ with the usual Euclidean inner product is a Riemannian manifold with constant curvature 00. $\mathfrak{g}_{x}$ can be seen as a generalization of inner products, where the norm of $p\in T_{x}\mathcal{M}$ is $\|p\|_{\mathfrak{g}}=\sqrt{\mathfrak{g}_{x}(p,p)}$. The choice of $\mathfrak{g}$ induces a global distance function $d(\cdot,\cdot)$ on $\mathcal{M}$. A geodesic between $x,y$ is a local distance minimizing smooth curve. In particular, the shortest paths are geodesics. With certain assumption on the structure of $\mathcal{M}$, one can define the exponential map $\exp_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ for $x\in\mathcal{M}$, and its inverse, the logarithmic map $\log_{x}:\mathcal{M}\to T_{x}\mathcal{M}$. Additionally, the parallel transport map $\mathrm{PT}_{x}(v,w)$, where $v,w\in T_{x}\mathcal{M}$, generalizes translation, transporting $w$ starting at $x$ in the direction of $v$ with no acceleration.
2.2 Deep Learning in Non-Euclidean Spaces
Recent years have witnessed an increasing interest in extending deep learning techniques to Riemannian manifolds. Here, we discuss several advances for designing neural networks and Transformers in non-Euclidean geometries, as well as optimization on manifolds. A more detailed discussion of recent advances can be found in Appendix A.
Non-Euclidean Neural Networks. Several works have explored neural networks that leverage geodesic distance to perform neural network operations (Bronstein et al., 2017; Masci et al., 2015; Boscaini et al., 2016) and in graph representation (Kong et al., 2022). Within hyperbolic learning, prior works have developed neural network layers (Ganea et al., 2018; Shimizu et al., 2020; Nickel & Kiela, 2017; Chen et al., 2021b; Weber et al., 2020), graph neural networks (Liu et al., 2019a; Chami et al., 2019), vision models (Bdeir et al., 2024; van Spengler et al., 2023), and residual neural networks (He et al., 2025), leveraging more effective embeddings for hierarchical structures. In addition, extensive works have developed equivariant neural networks that encode spherical geometry as inductive bias (Cohen et al., 2018; Esteves et al., 2017; Coors et al., 2018; Defferrard et al., 2020; Esteves et al., 2019). Neural networks for mixed curvature manifolds that encompass both hyperbolic and spherical models have also been proposed (Gu et al., 2019; Bachmann et al., 2020). Many Euclidean convex and stochastic optimization algorithms have been extended to manifold learning as well (Udriste, 1994; Zhang et al., 2016; Bécigneul & Ganea, 2018; Weber & Sra, 2021, 2022).
Non-Euclidean Transformers. Significant advancements have been made toward Transformers in non-Euclidean spaces in recent studies. Prior works have developed attention mechanisms and additional essential operations, such as layer normalization, to develop Transformers in hyperbolic, spherical, and mixed curvature manifolds (Gulcehre et al., 2019; Chen et al., 2021b; Shimizu et al., 2020; Yang et al., 2024c; Lai et al., 2023; Cho et al., 2023).
3 Foundation Models Should Embrace Non-Euclidean Geometries
Euclidean Foundation Models. Foundation models are typically trained on massive datasets to learn transferable representations that serve as a basis for downstream tasks (Bommasani et al., 2021). Transformer-based language models such as BERT (Devlin et al., 2018), GPT (Brown et al., 2020), and T5 (Raffel et al., 2020) are pre-trained on extensive text corpora. In computer vision, large-scale models such as Vision Transformer (ViT) (Dosovitskiy et al., 2020) and ResNet variants (He et al., 2016) enable state-of-the-art performance on tasks such as image classification and object detection. Multimodal foundation models like CLIP (Radford et al., 2021) and DALL-E (Ramesh et al., 2021) learn cross-domain embeddings for zero-shot image classification and text-to-image generation.
3.1 Limitations of Euclidean Geometry for Foundation Models
The Euclidean assumption is based on the idea that relationships between data points can be meaningfully characterized using distances measured in a flat, Euclidean space. However, theoretical and experimental verification has demonstrated that Euclidean geometry, with its isotropic nature and uniform scaling, fails to capture the complex structures of real-world data, resulting in significant distortions (Bourgain, 1986; Matoušek, 1996; Arias-de Reyna & Rodríguez-Piazza, 1992; Gupta, 1999; Matoušek, 1999). As a result, high-quality, low-distortion embeddings are often only possible in high-dimensional Euclidean space. Specifically, embeddings of complex structured data, such as hierarchies or trees, provably incur high rates of distortion (Bourgain, 1985; Matoušek, 1996). In this section, we highlight how the flat nature of the Euclidean space results in limitations and challenges for foundational models.
Non-Applicability of the Nash Embedding Theorem. The Nash Embedding Theorem states that any Riemannian manifold $\mathcal{M}$ of dimension $n$ admits an isometric embedding $f$ into $\mathbb{R}^{2n+1}$ (Nash, 1954), seemingly to imply that non-Euclidean spaces would only reduce the embedding dimension by half. However, the isometric embedding here is defined to preserve the Riemannian metric, meaning that it is locally distance preserving—the length of any path is preserved. However, for the shortest path between points $x,y\in\mathcal{M}$, its image under $f$ is not necessarily the shortest path (i.e., Euclidean straight line) between $f(x)$ and $f(y)$. Conversely, measuring the embedding distortion is concerned with whether a map is globally distance preserving. This occurs when the shortest path between $x$ and $y$ remains the shortest path between $f(x)$ and $f(y)$, which is defined by isometric embeddings between metric spaces. Note that an isometric embedding between Riemannian manifolds is in general not an isometric embedding between metric spaces.
In the context of foundational models, we are concerned with global distance-preserving embeddings, as computing the distance between any pairs of token embeddings is crucial for model training. As a result, the Nash Embedding Theorem is not applicable since global distortion could still arise from isometric embeddings between Riemannian manifolds. For this reason, by “isometry”, we refer to those between metric spaces, which captures the (bi-Lipschitz) distortion critical for foundational model embeddings. For more details on distortion and isometric embeddings, see Section A.2.
Dimensionality. High dimensionality is often required to embed complex structures in Euclidean space with low distortion. The following theorem demonstrates the distortion-dimension trade-off that Euclidean embeddings face even in the simple case of unweighted token relationships / self-attention, in the form of complete graphs.
Theorem 3.1.
(Matoušek (2002)) Let $X$ be an $n$-point metric space with uniform distance $1$, i.e., an unweighted complete graph with $n$ nodes. For $\epsilon>0$, the minimal $d$ such that $X$ can be embedded into $\mathbb{R}^{d}$ with distortion $(1+\epsilon)$ is $d=\Omega\left(\frac{\log(n)}{\epsilon^{2}\log(1/\epsilon)}\right)$
For any $p<2$, $\epsilon^{2}\log(1/\epsilon)$ tends to 00 at a faster rate than $\epsilon^{p}$ as $\epsilon\to 0$. As a result, Theorem 3.1 implies that $d$ grows near-quadratically with respect to inverse distortion, even when embedding simple structures. Furthermore, note that any un-weighted graph with $n$ nodes can be isometrically embedded into an unweighted complete graph with $n$ nodes. Thus Theorem 3.1 implies the same dimensionality issue for embedding any unweighted graph in Euclidean space.
Distortion. In the cases of more complex structures, non-trivial distortion could even exist regardless of the dimension of the Euclidean space due to its uniform scaling properties. The following theorem implies that a wide range of spaces cannot be isometrically embedded into Euclidean space, based on Markov convexity (Section A.3).
Theorem 3.2.
(Lee et al., 2007) Let $(X,d_{X}),(Y,d_{Y})$ be metric spaces. For every $p\in\mathbb{N}$, denote $\Pi_{p}(X),\Pi_{p}(Y)$ the Markov $p$-convexity constant of $X$ and $Y$ respectively. Let $c_{Y}(X)=\inf\{\mathrm{dist}(f):f:X\to Y\}$ denote the minimum distortion of embedding $X$ in $Y$. Then $c_{Y}(X)\geq\frac{\Pi_{p}(X)}{\Pi_{p}(Y)}$.
For instance, when $X$ models hierarchical token relationships, e.g., $X=B_{2^{k}}$ is a complete binary tree of depth $2^{k}$, we have $\Pi_{p}(X)\geq\sqrt{k}$ (Lee et al., 2007). Thus, the distortion for embedding binary trees of depth in any Euclidean space is at least $k$$c_{\mathbb{R}^{n}}(B_{k})\geq\Omega(1)\cdot\sqrt{\log k}$. Similarly, when $X$ represents circular or periodic dependencies in the tokens, e.g., $X$ is a ball of radius $r$ in a vertex-transitive graph, such as a cycle graph, the minimal distortion of embedding $X$ into $\mathbb{R}^{n}$ for any $n$ is $\Omega(\sqrt{\log r})$ (Lee et al., 2007).
Moreover, non-trivial distortion exists when embedding other forms of topological space as well, including the sphere $S^{k}\subseteq\mathbb{R}^{k+1}$, as shown in the following theorem.
Theorem 3.3.
(Robinson, 2006) Let $(X,d_{X})$ be a metric space with $X=\{a,b,c,d\}$ and $d_{X}(a,b)=d_{X}(a,c)=d_{X}(a,d)=2L$ and $d_{X}(b,d)=d_{X}(c,d)=L$ for $L\in\mathbb{R}^{+}$. Then $X$ admits no isometric embedding into $\mathbb{R}^{n}$ for any $n$.
As these points can be isometrically embedded into $S^{k}$, Theorem 3.3 shows that $S^{k}$ cannot be isometrically embedded into $\mathbb{R}^{n}$ for any $n\in\mathbb{N}$, resulting in distortion when encoding properties such as rotational equivariance. In contrast, non-Euclidean geometry (e.g., hyperbolic or spherical spaces) can provide a more natural representation of complex topological structures, reducing distortion and dimensionality of the embedding space. For instance, (Sarkar, 2011) showed that every finite tree admits an embedding into the hyperbolic plane $\mathbb{H}^{2}$ with only $1+\epsilon$ multiplicative distortion for any $\epsilon>0$. This leads to a distortion bound of $O(1)$ while requiring low dimensionality.
Table 1: Average (point-wise) distortion on canonical graphs with $96$ nodes, comparing four spaces with total dimension $6$. The least distortion is achieved by the space with the most suitable geometry.
Geometry Tree Cycle Ring of Trees
$|E|=95,|V|=96$ $|E|=96,|V|=96$ $|E|=96,|V|=96$
$\mathbb{R}^{6}$ $0.1036$ $0.1042$ $0.1060$
$\mathbb{H}^{-1,6}$ $\mathbf{0.0454}$ $0.2545$ $0.0736$
$\mathbb{S}^{1,6}$ $0.1440$ $\mathbf{0.0011}$ $0.1365$
$\mathbb{H}^{-1,3}\times\mathbb{S}^{1,3}$ $0.0624$ $0.1337$ $\mathbf{0.0686}$
Take-away. The implications of the previous theoretical discussion are numerous: (1) Limited scalability. Theorem 3.1 highlights the distortion-dimension trade-off for Euclidean foundation models when embedding complex structures, which is reflected in the computational resources required in these models. Non-Euclidean geometry produces higher quality embeddings in significantly lower dimensions, offering enhanced model scalability; (2) Performance bottleneck. Theorem 3.2 and 3.3 demonstrate that even in the case of an abundance of compute resources, the linear assumption in Euclidean foundation models could still incur significant distortion regardless of the embedding dimension for a wide range of topological structures, resulting in a performance upper bound.
Table 2: $\delta$-Hyperbolicity of the token embedding in various LLMs across several datasets. The bottom 2 rows show the $\delta$-hyperbolicity values of several metric spaces for reference.
Model arXiv C4 Common Crawl GitHub StackExchange Wikipedia
RoBERTa-Base (Liu et al., 2019b) $0.15\pm 0.06$ $0.18\pm 0.04$ $0.17\pm 0.04$ $0.12\pm 0.04$ $0.17\pm 0.07$ $0.07\pm 0.05$
LLaMA3.1-8B (Grattafiori et al., 2024) $0.15\pm 0.05$ $0.16\pm 0.07$ $0.15\pm 0.06$ $0.12\pm 0.05$ $0.18\pm 0.06$ $0.10\pm 0.04$
GPT-NeoX-20B (Black et al., 2022) $0.14\pm 0.03$ $0.17\pm 0.06$ $0.15\pm 0.05$ $0.11\pm 0.04$ $0.14\pm 0.04$ $0.09\pm 0.03$
Gemma2-9B (Team et al., 2024) $0.17\pm 0.06$ $0.19\pm 0.04$ $0.20\pm 0.05$ $0.15\pm 0.05$ $0.18\pm 0.04$ $0.15\pm 0.03$
Metric Space Sphere Space Dense Graph PubMed Graph Poincaré Space Tree Graph -
Reference $\delta$ values $0.99\pm 0.01$ $0.63\pm 0.01$ $0.40\pm 0.04$ $0.14\pm 0.01$ $0.0$
Figure 2: Token frequency v.s. token count (left 2) and token norm vs token count (right 2) for LLaMa3.1-8B and LLaMaGen. The datasets are chosen to be within the training corpus. The token-frequency figures show the scale-free properties of the token inputs. The token norms figures reflect this property for learned token embeddings to some extent, with token count decreasing exponentially for high normed tokens at the right tail. However, the Euclidean embeddings stil do not fully capture this property and deviate from it at the left tail. Statistics for additional models and dataset details are shown in Appendix B.
3.2 Non-Euclidean Geometry in Foundation Models
In this section, we empirically assess embedding distortions for different geometries to validate our claims in Section 3.1 and demonstrate that non-Euclidean geometry is more suitable. We then analyze token embeddings in foundation models, showing that structures that align with non-Euclidean geometry (Section 3.1) are prevalent, highlighting the need for alternative geometric frameworks.
Empirical Validation. We empirically validate our claim that Euclidean space fails to capture complex structures faithfully and that non-Euclidean spaces are better suited for producing high-quality embeddings. Table 1 compares the average (point-wise) distortion of four geometric spaces ($\mathbb{R}^{6}$, $\mathbb{H}^{-1,6}$, $\mathbb{S}^{1,6}$, and $\mathbb{H}^{-1,3}\times\mathbb{S}^{1,3}$) in representing three canonical graphs (Tree, Cycle, and Ring of Trees) with 96 nodes, each corresponding to a different type of intrinsic token relationships (hierarchical, cyclical, and both). The most suitable geometry varies by graph type—Lorentzian space ($\mathbb{H}^{-1,6}$) for trees, spherical space ($\mathbb{S}^{1,6}$) for cycles, and mixed geometry ($\mathbb{H}^{-1,3}\times\mathbb{S}^{1,3}$) for rings of trees—emphasizing the importance of selecting an appropriate geometry to minimize distortion.
Hierarchies in Token Embeddings. To explore the intrinsic structure within the representations of foundation models, we first utilize $\delta$-hyperbolicity (Gromov, 1987)—a curvature notion that quantifies the extent to which a metric space $(X,d)$ deviates globally from a perfect tree metric. Formal details are provided in Appendix C. Each token within a prompt is treated as a point in a discrete metric space $X$, and a graph is constructed based on similarity scores between each pair. We evaluate the hyperbolicity of token embeddings in LLMs, where lower values suggest a tree-like structure (Table 3). As shown in Table 2, the consistently low $\delta$-hyperbolicity values suggest hierarchical structures within each prompt across diverse datasets.
Additionally, we analyze the global token embedding distribution in LLMs and pre-trained vision models using datasets included in the models’ training corpus (Touvron et al., 2023; Grattafiori et al., 2024; Sun et al., 2024b). Figure 2 plots token input frequency distributions and their occurrences in the dataset on a log scale, revealing a scale-free structure among the token embeddings that follow power-law decay. The figure also shows token norm distributions for learned embeddings, where the count for high-norm embeddings decreases exponentially at the right tail, reinforcing the scale-free property. This scale-free organization suggests an underlying hierarchical structure (Barabási & Albert, 1999), where a small number of high-frequency tokens act as hubs within the semantic network. Such hierarchies are naturally represented in hyperbolic embedding spaces (Krioukov et al., 2010; Nickel & Kiela, 2017), which efficiently encode tree-like structures due to their exponential volume growth. The non-Euclidean structures in token distribution are exhibited to some extent even in Euclidean models and are most likely attributed to the models being optimized during training to maximize representational quality. However, the scale-free properties are still not yet fully captured by the Euclidean foundational model, where the count of embeddings with small norms still increases. See Appendix B for statistics of more models and the details of the datasets. We do observe that the non-Euclidean structure in the final embeddings is significantly reduced compared to earlier layers, even though some non-Euclidean structure persists in the final embeddings. Nevertheless, this does not diminish the presence of non-Euclidean structures in the model’s handling of inputs through token embeddings. After all, final embeddings focus on supporting prediction tasks after intermediate Euclidean transformations such as attention layers and layer normalization.
Additional Structures. In addition to hierarchical structure, data may exhibit other structural characteristics, such as cycles and loops. Foundation models built on Euclidean geometry struggle to represent such structures without significant distortion. Many real-world tasks, such as 3D shape analysis (Esteves et al., 2017, 2019), medical imaging (Bekkers et al., 2018; Winkels & Cohen, 2018), and physics-informed machine learning (Liao & Smidt, 2023; Liao et al., 2024; Anderson et al., 2019; Cohen et al., 2018), can benefit from encoding data geometry as inductive bias. Euclidean operations, such as the convolutional layers, encode only translation invariance (Esteves et al., 2020), resulting in performance limitations for these tasks.
3.3 The Necessity of Non-Euclidean Geometry for Foundation Models
Building on the previous findings, we further explore how encoding suitable geometric structure can improve foundation model performance.
(1) Addressing the limitations of Euclidean space in capturing intrinsic token structures. Recent research shows that the attention mechanism plays a pivotal role in the expressive capacity of LLMs (Aghajanyan et al., 2021; Song & Zhong, 2023; Vaswani et al., 2017; Balestriero et al., ).
Lemma 3.4 (Causal Single-Head Convex Hull (Balestriero et al., )).
Let $X\in\mathbb{R}^{T\times D^{(\ell)}}$ be the input to the $\ell$-th layer of an LLM, where $T$ is the sequence length and $D^{(\ell)}$ is the feature dimension. The output of attention head $h$ at position $i$ lies within the convex hull of the first $i$ rows of $XV_{h}^{(\ell)}$: $\text{Head}_{h}^{(\ell)}(X)_{i}\in\text{Hull}\left\{(V_{h}^{(\ell)})^{\top}x_{j}\mid j=1,\dots,i\right\}.$ Its effective dimension is bounded by the number of nonzero attention weights: $\text{dim}_{\text{eff}}\leq\#\left\{\text{Attn}_{h}^{(\ell)}(X)_{i,j}>0\mid j\in\{1,\dots,i\}\right\}.$ Here, $\text{Attn}_{h}^{(\ell)}(X)$ is the attention matrix for head $h$ at layer $\ell$, computed as: $\text{Attn}_{h}^{(\ell)}(X)=\text{softmax}_{\text{causal}}(XQ_{h}^{(\ell)}K_{h}^{(\ell)\top}X^{\top}).$
This lemma highlights that next-token prediction in LLMs is strongly influenced by relationships encoded in previous tokens. Standard attention mechanisms assume an Euclidean representation space, where token embeddings and their pairwise relationships fit this geometry. However, as investigated in Table 2, each prompt exhibits non-Euclidean characteristics. Consequently, the Euclidean attention mechanism fails to model hierarchical syntax, circular or periodic dependencies, and other complex relational structures without distortion, as demonstrated in Section 3.1. Replacing the Euclidean attention matrix with a non-Euclidean one (e.g., hyperbolic or spherical) could better capture previous token relationships, aligning with the intrinsic data structure and enhancing next-token prediction. For example, hyperbolic geometry compresses distances exponentially, ensuring that distant but structurally related tokens (e.g., a root concept and its distant occurrences in a prompt / document) remain meaningfully close. This enables attention mechanisms to efficiently preserve hierarchical relationships and capture long-range dependencies.
(2) Leveraging distortion-dimension trade-offs in hierarchical concept modeling. Recent studies extend beyond representation space analysis to unify it with the unembedding space, examining how Euclidean-based LLMs encode hierarchies geometrically (Park et al., 2024a, b). An LLM consists of a mapping function $\lambda$ and an unembedding layer. The function $\lambda$ maps input text $x$ to a vector $\lambda(x)\in\mathbb{R}^{d}$, while the unembedding layer assigns $\gamma(y)\in\mathbb{R}^{d}$ to each token $y$. The token probability distribution follows: $P(y\mid x)=\frac{\exp(\lambda(x)^{\top}\gamma(y))}{\sum_{y^{\prime}\in\text{Vocab}}\exp(\lambda(x)^{\top}\gamma(y^{\prime}))}.$ To unify the different spaces, the embedding and unembedding spaces can be reformulated using transformations $g(y)=A(\gamma(y)-\bar{\gamma}_{0}),\quad\ell(x)=A^{-\top}\lambda(x),$ where the Euclidean inner product serves as the causal inner product.
This framework, validated through experiments, reveals that Euclidean-based LLMs encode hierarchical concepts orthogonally, modeling a tree structure where parent (e.g., animal) and child (e.g., bird, mammal) vectors are perpendicular. However, this orthogonal representation demands high dimensionality, which scales significantly as hierarchical relationships grow more complex (Gupta, 1999). Non-Euclidean spaces offer a more efficient alternative, preserving hierarchical relationships while significantly reducing dimensionality (Nickel & Kiela, 2017, 2018).
(3) Surpassing Euclidean constraints in modeling multi-modal heterogeneity. Data from different modalities exhibit significant distributional variations, influenced by contextual factors, use cases, cultural differences, and varying interpretations of the same information. This complexity intensifies in multi-modal data, where each modality (e.g., images, text, audio) has distinct hierarchical structures (Li et al., 2022; Wei et al., 2023; Liang et al., 2022; Goel et al., 2022; Jiang et al., 2023). Many studies highlight this issue. Liang et al. (2022) showed that the latent modality gap and distinct modality structures emerge in the latent space due to initialization and the contrastive learning process, impacting downstream tasks. Wang et al. found that different modalities exist on separate manifolds, making a unified Euclidean foundation model highly redundant in parameters and requiring varying degrees of pruning for different modalities. This suggests that a unified Euclidean geometry struggles to capture multi-modal, cross-domain relationships, as its flat structure lacks the expressiveness needed for complex, multi-faceted interactions.
Hyperbolic geometry excels in vision-language foundation models by effectively capturing hierarchical relationships, enhancing representation interpretability (Desai et al., 2023; Pal et al., 2024; Ramasinghe et al., 2024). Its advantages extend to multimodal tasks, including image-video-skeleton (Li et al., 2024) and video-audio applications (Hong et al., 2023). For example, in hyperbolic embeddings, higher-level hierarchical concepts lie closer to the origin with more specific concepts residing in more peripheral regions, enabling geodesic reasoning when navigating through concept hierarchies. Hyperbolic embeddings minimize distortion while preserving real-world data structures, reinforcing the necessity of non-Euclidean geometry in foundation models. More broadly, generalized non-Euclidean geometries provide diverse curvatures, from positive to negative, enabling multiple manifolds to encode distinct data distributions (Wilson et al., 2014; Gao et al., 2022, 2021), surpassing the constraints of flat Euclidean space.
In summary, we argue that moving beyond Euclidean geometry is essential for next-generation foundation models. Euclidean models struggle with capturing token structures, face dimensionality limits, and fail to model multimodal heterogeneity, requiring high dimensions but lacking geometric adaptability. Non-Euclidean geometry, with its flexible curvature and manifold structures, overcomes these limitations in distortion, scalability, and adaptability, providing a more scalable and expressive framework for foundation models.
Figure 3: The roadmap for integrating non-Euclidean geometries into foundation models, includes (a) fine-tuning existing Euclidean foundation models, (b) pretraining from scratch, and (c) hybrid architectures. Four strategies are shown in (a), labeled with circled numbers 1-4, respectively: geometric prompt tuning, geometric low-rank adaptation, geometric knowledge distillation, and geometric transfer learning. All learnable components are highlighted in red in (a) and (c). We only show one decoder for each strategy for simplicity.
4 Towards Non-Euclidean Foundation Models
In the following roadmap on endowing foundation models with non-Euclidean geometry, we explore three progressive approaches: fine-tuning existing Euclidean models, building non-Euclidean models from scratch, and developing a hybrid framework combining both for optimal performance. We also discuss key implementation challenges of the roadmap, especially the computational efficiency of non-Euclidean operations.
4.1 Fine-tuning Existing Euclidean Foundation Models
The off-the-shelf Euclidean foundation model already encodes rich information from pretraining, offering a strong starting point for addressing the aforementioned challenges. A computationally efficient strategy is to adapt these models to non-Euclidean geometry, retaining their original capabilities while enabling generalization to data with non-Euclidean structures or relationships. We propose four strategies, shown in Figure 3(a): (1) Geometric prompt tuning; (2) Geometric low-rank adaptation; (3) Geometric knowledge distillation; and (4) Geometric transfer learning.
(1) Geometric Prompt Tuning. Prompt tuning offers a parameter-efficient alternative to full fine-tuning by introducing trainable, task-specific prompt tokens to the input. This approach helps mitigate catastrophic forgetting while requiring fewer trainable parameters. (Kirkpatrick et al., 2017; Han et al., 2024).
As there exist complex relations, such as hierarchical relationships between prefix prompts and downstream inputs, geometric prompts can be optimized through non-Euclidean spaces to align with the data geometry to better preserve structure and adapt to downstream tasks. For instance, trainable prompt and token embeddings could be introduced to better capture the topological relationships between prompts and text inputs.
(2) Geometric Low-Rank Adaptation. Low-Rank Adaptation (LoRA) offers an efficient way to adjust the model parameter space for downstream tasks given the observation of a low-rank structure (Hu et al., 2021). To equip the pre-trained model with non-Euclidean characteristics through geometric low-rank adaptation, we could operate the low-rank matrix multiplication directly on the manifold after projecting the input into non-Euclidean spaces, which avoids frequent switching and preserves the underlying geometric structure better (Yang et al., 2024a).
(3) Geometric Knowledge Distillation. Distilling knowledge into non-Euclidean spaces refers to transferring knowledge from a large, complex teacher model to a smaller, more efficient student model through geometric spaces. This approach utilizes the geometric properties of manifolds to teach the student to better inherit the manifold structure in the teacher model, such as by minimizing the gap between each layer’s output of both models, especially in high-dimensional spaces (Yang et al., 2022a; Hao et al., 2022) and resource-limited applications.
(4) Geometric Transfer Learning.
Geometric transfer learning aims to help foundation models learn across domains with aligned geometries, ensuring a much more effective and consistent knowledge transfer.
Geometry alignment objectives can be designed to supervise the transfer of geometric knowledge, such as hyperbolic contrastive learning for recommendation (Yang et al., 2024d; Ma et al., 2024), preserving the intrinsic structure of the target domain’s embedding space while retaining geometry-agnostic prior knowledge.
4.2 Pretraining from Scratch
Pretraining non-Euclidean foundation models requires addressing unique geometric challenges. We outline key components for adapting models to complex curvature-aware structures; see also Figure 3(b). A detailed mathematical formulation for each of the below sections
is presented in Table 4 in the Appendix.
Curvature Estimation. A data manifold’s curvature determines its intrinsic geometric properties, fundamentally affecting distance metrics, topological structure, and learning dynamics. Curvature estimation methods vary based on the data type. For graph data (e.g., networks, proteins), curvature is derived from topological properties, such as Ollivier-Ricci curvature or Gromov hyperbolicity (Ollivier, 2009; Jost & Liu, 2014; Yang et al., 2024a; Garcia Trillos & Weber, 2023). For non-graph data (e.g., text, images), curvature is typically estimated through learned embeddings (Khrulkov et al., 2020; Atigh et al., 2022) or techniques like Isomap (Tenenbaum et al., 2000) and UMAP (McInnes et al., 2018). We could also design learnable curvature within training pipelines using second-order statistics (Gao et al., 2022), reinforcement learning (Fu et al., 2021), and self-supervised learning (Sun et al., 2023, 2022).
Non-Euclidean Attention Mechanism. In non-Euclidean spaces, attention scores can be defined based on negative manifold distance $-d_{\mathcal{M}}(x,y)$ between queries and keys instead of dot products (Gulcehre et al., 2019; Shimizu et al., 2020; Chen et al., 2021b; Cho et al., 2023), with closer node pairs receiving higher attention weights. To aggregate attention, unified manifold centroids or tangent space operations can be used (Gu et al., 2019; Cho et al., 2023). Alternatively, linear attention mechanisms (Yang et al., 2024c) can be employed to improve computational efficiency by approximating traditional attention through unified tangent space operations.
Other Important Modules. Traditional Euclidean positional encoding, using sine and cosine functions (Vaswani et al., 2017), does not preserve the manifold structure in non-Euclidean spaces. Several approaches for non-Euclidean positional encoding include hyperbolic rotations (Yang et al., 2024c; Fein-Ashley et al., 2024), learned manifold encodings (Chen et al., 2021b), and direct embedding onto the manifold were proposed to represent token positions while maintaining geometric integrity. Residual connections should be formulated using isometric operations (He et al., 2025; van Spengler et al., 2023; Bdeir et al., 2024) to preserve geometric information across layers. Additionally, layer and batch normalization must be adapted to account for curvature (Yang et al., 2024c; Bdeir et al., 2024; van Spengler et al., 2023). One approach is to modify the softmax loss function and compute the probability distribution over tokens based on the manifold distance instead.
4.3 Hybrid Architectures
Hybrid architectures take a step further by merging both Euclidean and non-Euclidean foundation model architectures also provide a more universal inductive bias. We illustrate two promising strategies, also depicted in Figure 3(c).
Dynamic Geometry Adaptation Mechanisms. An intuitive way for hybrid modeling is to design a geometry-aware mechanism to shift dynamically between subspaces with different curvatures. The crux of this approach lies in how to perform efficient geometry adaption. A unified product manifold framework can serve as such a medium (Sun et al., 2024a), by utilizing a product layer to integrate diverse learnable curvature factors that adapt to different curvatures dynamically during the training process. Moreover, Mixture of Experts (MoE) (Zhou et al., 2022) provides a natural framework for such a hybrid paradigm using a geometry-aware sparse routing network aimed at minimizing average pairwise embedding distortion (Gu et al., 2019). This geometric routing selects the most appropriate Riemannian space considering data structure for the input (Guo et al., 2024), while avoiding interference from other spaces, thus naturally addressing issues of distortion and heterogeneity. Multi-Manifold Attention. To integrate more versatile underlying dependencies, we could transform self-attention into multi-manifold attention  (Konstantinidis et al., 2023; Kalitsios et al., 2024). Specifically, the input is embedded into a collection of manifolds (like Euclidean, hyperbolic, and spherical space) to represent differences in geometric structure across the data set. These geometric attention maps are then fused to produce a highly discriminative map for improved attention guidance.
4.4 Roadmap Implementation Challenges
Implementing and developing non-Euclidean foundation models come with unique challenges. One particular concern is that non-Euclidean operations are often complicated and require more computational resources, posing potential challenges when the scaling models to billions and trillions of parameters in practice. In particular, tangent-space-enabled methods and modules (Ganea et al., 2018; Chami et al., 2019; van Spengler et al., 2023), although intuitive by taking advantage of the Euclidean structure of the tangent spaces, incur significant computation overhead due to multiple mappings to and from the tangent bundle. In comparison, methods that operate directly on the manifold (Chen et al., 2021b; Yang et al., 2024c; He et al., 2025), while still computationally more expensive than Euclidean methods, typically have similar computational complexity as their Euclidean counterparts and offer significant speedup over tangent-space-based methods. This type of operation could be promising for managing the computational efficiency of non-Euclidean foundation models. Additionally, non-Euclidean models require fewer dimensions to embed complex structures when compared to Euclidean models, as seen in our discussion in Section 3.1. This enables the potential for non-Euclidean models to match the performance of Euclidean models with fewer parameters to build parameter-efficient foundation models. The reduction of parameter count could offset the computational overhead from non-Euclidean operations while offering additional benefits, such as the potential to continue the scaling law relationship between parameters and model performance.
Additional challenges include gaps in current research for more concrete analysis between embedding quality and model performance. Some previous works have explored how using manifolds that more accurately estimate the structure of the data can enhance performance in tasks related to GNNs and word embeddings (Gu et al., 2019). However, to the best of our knowledge, there currently lacks conclusive work directly connecting numerical distortion to downstream performance in foundation models, which is challenging as it could require prior knowledge of the ground-truth data geometry, compute resources to train multiple foundation models, and controlling and isolating the effects of distortion. Future works analyzing how distortion impacts the downstream tasks of foundation models would provide valuable insights to support better development of non-Euclidean methods for foundation models.
5 Alternative Views
While non-Euclidean geometries have clear theoretical benefits, there are practical challenges to consider. One of the main considerations is that operations such as parallel transport and exponential/logarithmic mappings add complexity, which might reduce some of the anticipated benefits in terms of efficiency. To tackle this, it is essential to develop specialized libraries that are optimized for these kinds of calculations, with efficient implementations of tensor operations that encode the underlying geometry.
Another view is that as hardware capacities increase, simply scaling Euclidean models to higher dimensions might reduce distortion. However, as pointed out by Theorem 3.2 and Theorem 3.3, there is an upper bound in how much distortion reduction is possible. Previous works have also empirically shown that non-Euclidean models outperform Euclidean models even with scaled parameter counts, such as in the case of equivariant and non-equivariant models (Brehmer et al., 2024). Additionally, in many domains, such as molecular structures or rare languages, data scarcity results in brute-force scaling being ineffective. Non-Euclidean geometries, on the other hand, can capture important relationships even in lower-dimensional settings (Sarkar, 2011), making them efficient in both parameter usage and data requirements, offering better performance scalability with respect to model size and a more reliable option for domains with limited high-quality data.
6 Conclusion
Foundation models benefit from embracing non-Euclidean geometry to resolve the inherent mismatch between Euclidean geometry and the non-Euclidean nature of real-world data. Non-Euclidean geometries reduce distortion for complex structures and relationships while enabling efficient representations, which is critical for trillion-parameter scaling. Aligning architectures with data geometry can potentially mitigate hallucinations, boost efficiency, and unlock heterogeneous scaling. We encourage the community to prioritize three directions: unified curvature-adaptive foundation model frameworks, geometry-aware benchmarks, and studying manifold-emergent capability links. Embracing this paradigm will catalyze AI systems that better reflect the rich geometries of human knowledge and physical reality.
Appendix A Comprehensive Background and Related Works
A.1 Riemannian Geometry and Non-Euclidean Foundations
Riemannian Manifolds. A smooth $n$-dimensional manifold $\mathcal{M}$ is a topological space in which each point $x\in\mathcal{M}$ has a neighborhood $U_{x}\subseteq\mathcal{M}$ that is locally Euclidean, meaning that there exists a homeomorphism between $U_{x}$ and an open subset of $\mathbb{R}^{n}$.
Tangent Space. Each point $x\in\mathcal{M}$ is associated with a tangent space $T_{x}\mathcal{M}$, which is an $n$-dimensional vector space serving as a first-order local approximation of $\mathcal{M}$ at $x$. This space encapsulates the possible directions in which one can move away from $x$ on the manifold.
Riemannian Metric. A Riemannian metric $\mathfrak{g}$ on $\mathcal{M}$ is a collection of positive-definite bilinear forms $\mathfrak{g}_{x}(\cdot,\cdot):T_{x}\mathcal{M}\times T_{x}\mathcal{M}\to\mathbb{R}$, smoothly varying with $x$. The metric $\mathfrak{g}_{x}$ induces the sectional curvature at each point, which measures the extent to which the manifold deviates from flatness at $x$. A Riemannian manifold is then defined as the pair $(\mathcal{M},\mathfrak{g})$. For instance, $\mathbb{R}^{n}$ with the usual Euclidean inner product is a Riemannian manifold with zero curvature. The metric $\mathfrak{g}_{x}$ generalizes the notion of inner products, with the norm of a vector $p\in T_{x}\mathcal{M}$ given by $\|p\|_{\mathfrak{g}}=\sqrt{\mathfrak{g}_{x}(p,p)}$. The choice of the Riemannian metric also induces a global distance function $d(\cdot,\cdot)$ on $\mathcal{M}$.
Geodesic. A geodesic between two points $x$ and $y$ is a smooth curve that locally minimizes the distance between these points. In particular, the shortest path between $x$ and $y$ is a geodesic.
Exponential Map. Under certain conditions, one can define the exponential map $\exp_{x}:T_{x}\mathcal{M}\to\mathcal{M}$, which lifts points from the tangent space $T_{x}\mathcal{M}$ to the manifold $\mathcal{M}$, by associating a vector in $T_{x}\mathcal{M}$ to a point on $\mathcal{M}$ along a geodesic.
Logarithmic Map. The logarithmic map $\log_{x}:\mathcal{M}\to T_{x}\mathcal{M}$ is the inverse of the exponential map, provided certain assumptions on $\mathcal{M}$ hold.
Geodesics and Geodesic Operations. The Riemannian metric $\mathfrak{g}_{x}$ can be viewed as a generalization of the inner product, where the norm of a vector $p\in T_{x}\mathcal{M}$ is defined by $\|p\|_{\mathfrak{g}}=\sqrt{\mathfrak{g}_{x}(p,p)}$. The choice of $\mathfrak{g}$ induces a global distance function $d(\cdot,\cdot)$ on $\mathcal{M}$, where geodesics are the locally distance-minimizing curves. The length of a geodesic between two points determines the geodesic distance. The exponential map $\exp_{x}$ maps a vector $v\in T_{x}\mathcal{M}$ to a point on $\mathcal{M}$ along the geodesic starting at $x$. The logarithmic map $\log_{x}$ is the inverse of this process. Additionally, the parallel transport map $\mathrm{PT}_{x}(v,w)$ transports vectors along geodesics, providing a canonical way to move vectors in a manner consistent with the underlying geometric structure. It canonically transports a vector $w$ along a geodesic emanating from $x$ with initial velocity $v$ and zero acceleration. This generalizes the classical notion of translation in Euclidean space.
Hyperbolic Spaces. Hyperbolic spaces are Riemannian manifolds with constant negative curvature, i.e., with curvature $-K<0$. Common models for hyperbolic space include the Poincaré ball model $\mathbb{P}^{K,n}$ and the Lorentz hyperboloid $\mathbb{L}^{K,n}$, which have been extensively studied in the context of deep learning (Nickel & Kiela, 2017; Ganea et al., 2018). For points $\mathbf{x},\mathbf{y}\in\mathbb{L}^{K,n}$, their inner product $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}$ is given by $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=-x_{t}y_{t}+\mathbf{x}_{s}^{T}\mathbf{y}_{s}=\mathbf{x}^{T}\mathfrak{g}_{n}^{K}\mathbf{y}$ with $|\|\mathbf{x}\||_{\mathcal{L}}\coloneq\sqrt{|\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}|}$ being the Lorentzian norm. Formally, $\mathcal{L}^{n}$ is the point set $\mathcal{L}^{n}=\{\mathbf{x}\in\mathbb{R}^{n+1}:\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=1/K,x_{t}>0\}.$ $\mathbb{P}^{n,K}$ is the $n$-dimensional sphere $S^{n}$ with radius $1/\sqrt{K}$, with the Riemannian metric $g_{x}^{\mathbb{P}}=\lambda_{x}^{2}g^{E}$, where $\lambda_{x}:=\frac{2}{1-c\|x\|^{2}}$ and $\mathfrak{g}^{E}$ is the Euclidean metric. Other models, such as the Klein model, also exist. These models are isometric, meaning that there is a smooth correspondence between points in different models that preserves distances, angles, and geodesics. This property allows for the selection of the most suitable model for a given application.
Spherical Spaces. Spherical spaces are Riemannian manifolds with constant positive curvature, i.e., with curvature $K>0$. An $n$-dimensional spherical space $\mathbb{S}^{K,n}$ is an $n$-dimensional sphere of radius $K^{-\frac{1}{2}}$, equipped with the Riemannian metric induced by the Euclidean metric on $\mathbb{R}^{n+1}$.
Mixed Curvature Spaces. A mixed curvature space $\mathcal{M}$ is defined as a product manifold consisting of Euclidean, spherical, and hyperbolic spaces. The Riemannian metric and geodesic operations for such a manifold are defined component-wise, enabling effective computational implementation for downstream tasks.
Generalized Riemannian Manifolds. Generalizations of Riemannian manifolds can be obtained by relaxing some of the assumptions in their classical definition. One notable generalization is the pseudo-Riemannian manifold, in which the metric $\mathfrak{g}$ is an indefinite bilinear form, allowing for both positive and negative signs. This generalization is useful in contexts such as relativistic physics, where spacetime is modeled as a pseudo-Riemannian manifold.
A.2 Non-Applicability of the Nash Embedding Theorem
The Nash Embedding Theorem roughly states that any Riemannian manifold of dimension $n$ admits an isometric embedding into $\mathbb{R}^{2n+1}$ (Nash, 1954). While it may appear as if this allows for Euclidean embeddings of complex structures with no distortion and only twice the dimension, this is in fact a confusion in vocabulary between the notion of isometric embeddings between those of Riemannian manifolds and those of metric spaces.
Definition A.1.
Let $(\mathcal{M},\mathfrak{g}),(\mathcal{M}^{\prime},\mathfrak{g}^{\prime})$ be Riemannian manifolds. An isometric embedding of Riemannian manifolds is a smooth map $f:\mathcal{M}\to\mathcal{M^{\prime}}$ such that $\mathfrak{g}=f^{*}\mathfrak{g^{\prime}}$. Let $(X,d_{X}),(Y,d_{Y})$ be metrics spaces. An isometric embedding of metric spaces is a map $f:X\to Y$ such that $d_{X}(a,b)=d_{Y}(f(a),f(b))$ for all $a,b\in X$.
Hence in the former, which is also the isometric embedding afforded by the Nash Embedding Theorem, the map $f$ preserves the Riemannian metric, i.e. the inner product on the tangent bundle. As a result, the isometry is locally distance preserving, in the sense that length of any path is preserved under $f$. However, given points $x,y$ connected by a shortest path $\gamma$, the straight line path connecting $f(x),f(y)$ in the co-domain is not necessarily $f(\gamma)$ (note that $f$ need not to be surjective). As a result, measuring the distortion of embeddings is concerned with whether $f$ is globally distance preserving, or whether the shortest distance between $f(x)$ and $f(y)$ is the length of $f(\gamma)$, which is defined by isometric embeddings between metric spaces. Note that an isometric embedding of Riemannian manifolds is in general not an isometric embedding of metric spaces. For instance, given the sphere $S^{1}$, its usual Riemannian metric is inherited from the Riemannian metric for $\mathbb{R}^{2}$, i.e. the usual inner product. The identity map is then an isometric embedding $S^{1}\hookrightarrow\mathbb{R}^{2}$ as Riemannian manifolds. However, the distance between points on the sphere does not coincide with the Euclidean distance of their image. As an example, antipodal points have distance $\pi$ in $S^{1}$ but distance $1$ in $\mathbb{R}^{2}$.
In the context of foundational models, we are concerned with globally distance preserving embeddings, as computing the distance between any pairs of token embeddings is crucial for model training. As a result, the Nash Embedding Theorem is not applicable since global distortion could still arise from isometries between Riemannian manifolds. For this reason, by ”isometry”, we refer to those between metric spaces unless otherwise specified, which captures the notion of distortion critical for foundational model embeddings.
Definition A.2.
Let $(X,d_{X}),(Y,d_{Y})$ be metric spaces equipped with the respective distance metrics and $f:X\to Y$ be a map. The bi-Lipschitz distortion of $f$ is $\mathrm{dist}(f)=\|f\|_{\mathrm{Lip}}\|f^{-1}\|_{\mathrm{Lip}}$, where $\|f\|_{\mathrm{Lip}}$ is the (possibly infinite) Lipschitz-constant of $f$. For a pair of points $(a,b)\in X^{2}$, the point-wise distortion is given by $\frac{|d_{X}(a,b)-d_{Y}(f(a)-f(b))|}{d_{X}(a,b)}$.
Both notions of distortion measure the deviation of $f$ from an isometry between metric spaces. Note that the minimum distortion in the case of bi-Lipschitz distortion is $1$.
A.3 Markov Convexity
In this section we provide the relevant background on notion of Markov convexity. Let $(X,d_{X})$ be a metric space. Then the Markov $p$–convexity constant $\Pi$ (for a fixed positive integer $p$) of the metric space $X$ is a universal constant (or $\infty$) define as follows:
Definition A.3.
$\Pi$ is the smallest constant s.t. for any Markov chain on $(X_{t})_{t\geq 0}$ on a state space $\Omega$, and every map $f:\Omega\to X$, and for any $m\in\mathbb{N}$, we have
$\sum_{n=0}^{\infty}\frac{1}{2^{np}}\sum_{t\in\mathbb{Z}}\mathbb{E}\!\Bigl{[}d\bigl{(}f(X_{t}),f(X_{t+2^{n}})\bigr{)}^{p}\Bigr{]}\;\;\leq\;\;\Pi^{p}\sum_{t\in\mathbb{Z}}\mathbb{E}\!\Bigl{[}d\bigl{(}f(X_{t}),f(X_{t+1})\bigr{)}^{p}\Bigr{]}$
Roughly speaking, when $\Pi<\infty$, the $p$-th moment of one-step increments dominates the $p$-th moments of exponential length steps. Intuitively, measures how tightly local behaviors in $X$ control and estimate global behaviors on the space, with lower values $\Pi$ showing tighter control.
A.4 Non-Euclidean Structure in the Real World
Non-Euclidean Structures in Natural Language Processing. Language exhibits inherently hierarchical structures - from concept taxonomies to entailment relationships - that challenge traditional Euclidean representations. These hierarchical relationships between linguistic units naturally manifest on non-Euclidean manifolds, particularly in hyperbolic space, which has emerged as a powerful framework for natural language processing (Dhingra et al., 2018; Leimeister & Wilson, 2018). Foundational work has demonstrated that hyperbolic embeddings can effectively capture word-level semantics (Tifrea et al., 2019) and concept hierarchies (Le et al., 2019), leveraging the exponential volume growth of hyperbolic space to model tree-like linguistic structures. The success of hyperbolic representations has sparked various advanced applications: from question answering systems (Tay et al., 2018), privacy-preserving text representations (Feyisetan et al., 2019), to multi-document summarization that captures document-level discourse structure (Song et al., 2023). Recent work has further extended these approaches to cross-lingual settings (Saxena et al., 2022) and contextual language models (Chen et al., 2021a), demonstrating the broad utility of non-Euclidean geometries in modern natural language processing.
Non-Euclidean Structures in Computer Vision. Similar to NLP, many computer vision tasks involve data that naturally resides in intricate manifolds that are challenging to model using conventional Euclidean space (Mettes et al., 2024). For instance, visual entities often form inherent hierarchical relationships among object classes, between scenes and their constituent categories (Ge et al., 2023; Pal et al., 2024), or scenes at varying levels of granularity (Khrulkov et al., 2020). In these scenarios, hyperbolic geometry provides a compelling alternative to the Euclidean representations in representing the exponential growth of hierarchical structures with minimal distortion (Sala et al., 2018). Its advantages have been demonstrated across a wide range of applications, including image segmentation (Atigh et al., 2022), action classification (Chen et al., 2022a) video prediction (Surís et al., 2021), deformable 3D surfaces (Masci et al., 2015). In parallel, hyperspherical learning has become integral to modern contrastive learning with cosine similarity, underpining tasks ranging from self-supervised learning (Durrant & Leontidis, 2022) to long-tailed classification (Kasarla et al., 2022) and few-shot learning (Trosten et al., 2023).
Non-Euclidean Structures in Complex Networks. Networks, whether they represent social interactions, user purchasing preferences, or transportation systems, often exhibit complex, non-Euclidean relationships that traditional Euclidean models fail to capture effectively. Social networks, for example, are best described by graph structures where nodes (individuals) are connected by edges (relationships) that can be directional, weighted, or even exhibit hierarchical properties. These networks typically involve intricate dependencies and nonlinear relationships, requiring geometric frameworks beyond Euclidean space to model effectively.
Non-Euclidean Structures in Natural Sciences. In natural science, many systems exhibit intricate structures that Euclidean space struggles to capture effectively. In biology, non-Euclidean geometries are integral to analyzing and modeling complex organic structures, such as protein folding (Villegas-Morcillo et al., 2021), single-cell RNA-seq data (Klimovskaia et al., 2020; Ding & Regev, 2021; Bhasker et al., 2024), and phylogenetic trees (Matsumoto et al., 2021), where hyperbolic and spherical geometries are commonly observed. In neuroscience, hyperbolic geometry is shown to be more effective than Euclidean counterpart in modeling the brain’s cortical folding (Urdapilleta et al., 2015), brain surface (Shi et al., 2013), and hippocampal spatial representations (Zhang et al., 2023), aiding in the study of spatial organization and connectivity.
A.5 Deep Learning with Non-Euclidean Geometries
Recent years have witnessed an increasing interest in extending deep learning techniques to Riemannian manifolds. Here we discuss in further detail the advances for designing neural networks and Transformers in non-Euclidean geometries, as well as optimization on manifolds.
Geodesic Neural Networks. Geodesic neural networks leverage geodesic, particularly geodesic distances, to perform neural operations that preserve geometric structure on manifold-structured data (Bronstein et al., 2017). Several works have developed geodesic convolutional layers by applying filters to local patches in geodesic polar coordinates (Masci et al., 2015), learning directionally sensitive filters along principal curvature directions (Boscaini et al., 2016), or learnable kernel functions that operate on local coordinate systems (Monti et al., 2017). More recent works such as GDGNN (Kong et al., 2022) have integrated geodesic operations with graph representations.
Hyperbolic Neural Networks. Hyperbolic neural networks exploit the geometry of hyperbolic space to learn embeddings that reflect hierarchical relationships more effectively than their Euclidean counterparts (Nickel & Kiela, 2017). HNN (Ganea et al., 2018) and HNN++(Shimizu et al., 2020) developed many basic operations, such as hyperbolic linear and convolutional layers, and multinomial logistic regression (MLR). HGCN (Chami et al., 2019) and HGNN (Liu et al., 2019a) were then among the first to develop hyperbolic graph neural networks (GNNs). More recently, HyboNet (Chen et al., 2021b) proposed a framework of hyperbolic neural networks that does not depend on the Euclidean tangent spaces; Poincaré ResNet (van Spengler et al., 2023) and HCNN (Bdeir et al., 2024) developed components for hyperbolic vision models; LResNet (He et al., 2025) proposed an efficient and stable residual connection method.
Spherical Neural Networks. Spherical neural networks are designed for data that naturally reside on spheres or benefit from spherical symmetry. Spherical CNNs (Cohen et al., 2018; Esteves et al., 2017) extended convolutions and pooling to preserve rotational symmetries. SphereNet (Coors et al., 2018) introduced a framework for learning spherical image representations by encoding distortion invariance into convolutional filters. DeepSphere (Defferrard et al., 2020) proposed a graph-based approach. SWSCNN (Esteves et al., 2020) later proposed a fully spherical CNN that allows for anisotropic filters.
Mix-curvature Neural Networks. Mix-curvature neural networks uses product spaces of the aforementioned manifolds to better model data that have local neighborhoods exhibiting different geometric properties.  (Gu et al., 2019) developed the first learning framework on product spaces, introducing fundamental techniques such as mean and loss functions for embedding optimization. $\kappa$-GCN (Bachmann et al., 2020) then extended learning on product spaces to GCNs, introducing a unified and differentiable Gyrovector spaces framework to constant curvature spaces beyond hyperbolic manifolds.
Non-Euclidean Transformers. Significant advancements have been made toward Transformers in non-Euclidean spaces in recent studies. Within hyperbolic learning, several works have proposed hyperbolic self-attention mechanisms (Gulcehre et al., 2019; Chen et al., 2021b; Shimizu et al., 2020) and hyperbolic linear attentions (Yang et al., 2024c), enabling constructions of hyperbolic Transformers. Hyperbolic fine-tuning methods have also been developed for LLMs (Yang et al., 2024b). Recent works have also proposed hyperbolic vision Transformers (Fein-Ashley et al., 2024). Attention mechanisms have been developed for spherical spaces as well (Lai et al., 2023). Further, Transformers have been developed for mixed curvature manifolds as well (Cho et al., 2023).
Optimization on manifolds. Learning on manifolds often times require optimizing parameters with manifold constraints. Many classical convex optimization algorithms have been extended to the manifold-valued setting (Udriste, 1994; Bacák, 2014; Zhang & Sra, 2016; Weber & Sra, 2022). Stochastic optimization on manifolds has been studied extensivley (Bonnabel, 2013; Zhang et al., 2016; Bécigneul & Ganea, 2018; Weber & Sra, 2021), which includes extensions of algorithms such as SGD and Adam, which are suitable for training models on geometric domains.
Appendix B Additional Statistics and Dataset Details
In this section we give details regarding the datasets we used, as well as the show more statistic results for more LLMs.
B.1 Dataset Details
For the evaluation of token embedding distribution in LLMs, we incorporated a wide range of datasets, including a subset of the RedPajama dataset (Weber et al., 2024) encompassing the arXiv, C4, Common Crawl, GitHub, Wikipedia, and StackExchange datasets; math reasoning datasets such as GSM8K (Cobbe et al., 2021), MATH50K (Hendrycks et al., 2021),MAWPS (Koncel-Kedziorski et al., 2016), and SVAMP (Patel et al., 2021); and common sense reasoning datasets, including BoolQ, WinoGrande (Sakaguchi et al., 2019), and OpenBookQA (Mihaylov et al., 2018).
B.2 More Statistics
In Figure 4 we show the statistics for token embeddings for more LLMs, including GPT-NeoX-20B (Black et al., 2022), OPT-13B (Zhang et al., 2022), RoBERT-Base (Liu et al., 2019b), Gemma2-9B (Team et al., 2024), LLaMa3.1-8B (Grattafiori et al., 2024), and LLaMa-13B (Touvron et al., 2023). The top 2 rows show distribution of the norm of the token embeddings and the bottom 2 rows show the distribution of the frequency of each token embedding. The token frequency distribution demonstrate scale-free property with power law decay, whereas the token norm show rapid decreases in token count for higher normed tokens at the right tail. However, still none of the Euclidean foundational models fully capture the underlying scale-free property of the distribution, with all of them having an initial increase in token count against token norm for small normed token embeddings.
Figure 4: Token embeddings statistics for GPT-NeoX-20B, OPT-13B, RoBERT-Base, Gemma2-9B, LLaMa3.1-8B, and LLaMa-13B. The top 2 rows show distribution of the norm of the token embeddings and the bottom 2 rows show the distribution of the frequency of each token embedding. The token frequency distribution demonstrate scale-free property with power law decay, whereas the token norm show rapid decreases in token count for higher normed tokens at the right tail.
Appendix C $\delta$-Hyperbolicity Computation
Given any four points $a$, $b$, $c$, and $w$ in a metric space, the Gromov product $[a,c]_{w}$ at $w$ is bounded below by the minimum of the Gromov products $[a,b]_{w}$ and $[b,c]_{w}$, minus a slack term $\delta$:
$[a,c]_{w}\geq\min([a,b]_{w},[b,c]_{w})-\delta.$ (1)
The Gromov product between $a$ and $b$ with respect to $w$ is defined as:
$[a,b]_{w}=\frac{1}{2}\left(d(a,w)+d(b,w)-d(a,b)\right).$ (2)
A metric space $X$ is said to be $\delta$-hyperbolic if this inequality holds for all choices of $a$, $b$, $c$, and $w$. In geodesic metric spaces, $\delta$-hyperbolicity implies that geodesic triangles satisfy the $\delta$-slim property, meaning that any point on one side of a geodesic triangle is at most a distance of $\delta$ from some point on one of the other two sides.
In an exact tree metric, where the sides of any triangle intersect at a single point, the hyperbolicity constant $\delta$ is zero. This follows from the fact that the four-point condition holds as an equality for all points in the space.
Table 3: Hyperbolicity values $\delta$ for different metric spaces.
Sphere Space Dense Graph PubMed Graph Poincare Space Tree Graph
$\delta$ $0.99\pm 0.01$ $0.62\pm 0.01$ $0.40\pm 0.04$ $0.14\pm 0.01$ $0.0$
Appendix D Foundational Operations for Pretraining Non-Euclidean Foundation Models
Table 4: Geometric Foundation Model Operations: Euclidean vs. Manifold Formulations
Operation Euclidean Space Manifold Space
Curvature (K) $K=0$ $K\in\mathbb{R}$
Attention Score $\alpha_{qk}=\textrm{softmax}\left(\frac{q\cdot k^{\top}}{\sqrt{d_{k}}}\right)$ $\alpha_{qk}^{\mathcal{M}}=\textrm{softmax}\left(\frac{-d_{\mathcal{M}}^{2}(q,k)}{\sqrt{d_{k}}}\right)$
Rotary PE $Q_{i}^{\text{RoPE}}=Q_{i}\operatorname{Rot}(\mathbf{p}_{i})$; $K_{i}^{\text{RoPE}}=K_{i}\operatorname{Rot}(\mathbf{p}_{i})$ $Q_{i}^{\text{RoPE}_{\mathcal{M}}}=\operatorname{PT}_{\mathcal{M}}(Q_{i},\mathbf{p}_{i})$; $K_{i}^{\text{RoPE}_{\mathcal{M}}}=\operatorname{PT}_{\mathcal{M}}(K_{i},\mathbf{p}_{i})$
Residual Connection $x^{(l+1)}=x^{(l)}+f(x^{(l)})$ $x^{(l+1)}=\exp_{x^{(l)}}(\lambda\cdot f(x^{(l)}))$
Layer Norm $\hat{x}=\frac{x-\mu}{\sigma}$ $\hat{x}=\exp_{\mu_{\mathcal{M}}}\left(\frac{\log_{\mu_{\mathcal{M}}}(x)}{\sigma_{\mathcal{M}}}\right)$
Cross-Entropy Loss $\mathcal{L}=-\sum_{t}\log p_{t}$ $\mathcal{L}=-\sum_{t}\log\frac{\exp(-d_{\mathcal{M}}(z_{t},z^{*}))}{\sum_{t^{\prime}}\exp(-d_{\mathcal{M}}^{2}(z_{t},z_{t^{\prime}}))}$
Optimization $\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}J(\theta)$ $\theta_{t+1}=\operatorname{Ret}_{\theta_{t}}\left(-\eta\text{Proj}_{T_{\theta_{t}}\mathcal{M}}\nabla J\right)$
FFN $y=W_{2}\sigma(W_{1}x+b_{1})+b_{2}$ $y=\exp_{\mathbf{0}}\left(W_{2}\sigma(\log_{\mathbf{0}}(W_{1}\otimes x\oplus b_{1}))\right)$
Attention Aggregation $h=\sum_{i}\alpha_{i}v_{i}$ $h=\operatorname{WeightedExpSum}(\{v_{i}\},\{\alpha_{i}\})$
  • $\operatorname{PT}_{\mathcal{M}}$:Parallel transport preserving vector properties during translation; $\exp_{\mu_{\mathcal{M}}}$: Exponential map from tangent space at Fréchet mean $\mu_{\mathcal{M}}$; $\log_{\mu_{\mathcal{M}}}$: Inverse exponential map projecting to tangent space; $\operatorname{Ret}$: Retraction mapping for parameter updates; Proj: Tangent space projection operator
Table 4 systematically compares foundational operations in Euclidean space with their adaptations to non-Euclidean manifold spaces, highlighting critical geometric modifications required for pretraining curvature-aware foundation models. Below, we explain the key components and their mathematical formulations:
Curvature (K). In Euclidean space, curvature is fixed at $K=0$, reflecting flat geometry. In manifold spaces, curvature $K\in\mathbb{R}$ is a learnable or estimated parameter that defines the intrinsic geometry (hyperbolic $K<0$, spherical $K>0$, or mixed). This value influences all subsequent operations, requiring dynamic adjustments to distance metrics and parameter updates. Curvature estimation methods (e.g., Ollivier-Ricci for graphs or learned embeddings for non-graph data) ensure geometric consistency across tasks.
Attention Mechanism. Euclidean attention computes similarity via dot products $\alpha_{qk}=\textrm{softmax}\left(\frac{q\cdot k^{\top}}{\sqrt{d_{k}}}\right)$, while manifold attention replaces this with geodesic distance: $\alpha_{qk}^{\mathcal{M}}=\textrm{softmax}\left(\frac{-d_{\mathcal{M}}^{2}(q,k)}{\sqrt{d_{k}}}\right)$. The negative squared distance prioritizes proximity on the manifold, preserving geometric relevance. Aggregation uses weighted Fréchet means (via exponential maps) or tangent space projections to combine features without violating curvature constraints.
Positional Encoding (Rotary PE). Euclidean positional encodings apply rotation matrices $\operatorname{Rot}(\mathbf{p}i)$ to query/key vectors. For manifolds, parallel transport $\operatorname{PT}_{\mathcal{M}}$ replaces rotations, translating positional shifts along geodesics while preserving vector orientation relative to the manifold’s curvature. This ensures positional relationships respect intrinsic geometry.
Residual Connections. Standard residuals $x^{(l+1)}=x^{(l)}+f(x^{(l)})$ are replaced by manifold equivalents: $x^{(l+1)}=\exp_{x^{(l)}}(\lambda\cdot f(x^{(l)}))$. Here, the exponential map $\exp$ projects tangent space updates $f(x^{(l)})$ onto the manifold, scaled by $\lambda$, to preserve geometric stability across layers.
Layer Normalization. Euclidean layer norm standardizes features via $\hat{x}=\frac{x-\mu}{\sigma}$. On manifolds, operations occur in the tangent space at the Fréchet mean $\mu_{\mathcal{M}}$: $\hat{x}=\exp_{\mu_{\mathcal{M}}}\left(\frac{\log_{\mu_{\mathcal{M}}}(x)}{\sigma_{\mathcal{M}}}\right)$, where $\log_{\mu_{\mathcal{M}}}$ maps points to the tangent space for normalization before reprojection.
Cross-Entropy Loss. The manifold loss $\mathcal{L}=-\sum_{t}\log\frac{\exp(-d_{\mathcal{M}}(z_{t},z^{*}))}{\sum_{t^{\prime}}\exp(-d_{\mathcal{M}}^{2}(z_{t},z_{t^{\prime}}))}$ replaces Euclidean dot products with geodesic distances, ensuring probabilities reflect the manifold’s geometry. This penalizes deviations in the curved space rather than in a flat embedding.
Optimization. Euclidean SGD $\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}J(\theta)$ is adapted via retractions $\operatorname{Ret}_{\theta_{t}}$, which map gradient steps $-\eta\text{Proj}_{T_{\theta_{t}}\mathcal{M}}\nabla J$ from the tangent space back to the manifold, ensuring updates respect curvature constraints.
Feed-Forward Network (FFN). Manifold FFNs $y=\exp_{\mathbf{0}}\left(W_{2}\sigma(\log_{\mathbf{0}}(W_{1}\otimes x\oplus b_{1}))\right)$ use Möbius operations ($\otimes$,$\oplus$ , ⊕) for linear transformations and biases, followed by activation in the tangent space. The exponential map $\exp_{\mathbf{0}}$ ensures outputs remain on the manifold.
Attention Aggregation. Instead of weighted sums $h=\sum_{i}\alpha_{i}v_{i}$, manifolds use $\operatorname{WeightedExpSum}$, which computes Fréchet means of values $v_{i}$ weighted by $\alpha_{i}$, ensuring aggregated features lie on the manifold.
These adaptations collectively enable pretraining in non-Euclidean spaces by preserving geometric integrity. Operations like parallel transport, exponential/log maps, and retractions ensure compatibility with curvature, while specialized normalization and loss functions align learning dynamics with the manifold’s intrinsic structure. The table underscores the necessity of redefining core components—from attention to optimization—to build effective foundation models for hyperbolic and mixed-curvature geometries.